OpenStudy (anonymous):

If I have a 2 x matrix to the nth power. What does that mean. I mean let's say n equals 2.

OpenStudy (jamesj):

such as \[ A^2 \]? It means exactly what you think it means. Multiply the matrix by itself 2 times. \[ A^2 = A A \] \[ A^3 = A A A \] etc.

OpenStudy (mr.math):

Lets say you have a matrix A, then \(A^2\) mean \(A\timesA\).

OpenStudy (mr.math):

\(A\times A\).

OpenStudy (jamesj):

notice that raising a matrix to an exponent only makes sense of the matrix is square. Otherwise you can't do the matrix multiplication.